OpenStudy (anonymous):

Exact value of sec(-11pi/12)

OpenStudy (jdoe0001):

\(\bf -\cfrac{11}{12}\implies \cfrac{9}{12}-\cfrac{20}{12}\implies \cfrac{3}{4}-\cfrac{5}{3}\qquad thus \\ \quad \\ -\cfrac{11\pi}{12}\implies \cfrac{3\pi}{4}-\cfrac{5\pi}{3}\qquad thus \\ \quad \\ sec\left(-\frac{11\pi}{12}\right)\implies sec\left(\frac{3\pi}{4}-\frac{5\pi}{3}\right)\implies \cfrac{1}{cos\left(\frac{3\pi}{4}-\frac{5\pi}{3}\right)}\) and you can get those values from your Unit Circle

OpenStudy (anonymous):

thank you ^.^

OpenStudy (jdoe0001):

yw